A transverse wave described by y = (0.02 m) sin [(1.0 m–1)x + (30 s–1)] propagates on a stretched string having a linear mass density of 1.2 x 10–4 kg/m. Find the tension in the string.

Grade:12

1 Answers

S sai Suhrut
6 years ago
let T= tension 
linear masss density (m) =1.2x 10^-4 kg/m
we know wave speed = w/k= 30/1=30m/s
imples that 
we know   v=√(T/m)
implies that
30=√(T/m)
=> 900* 1.2 x 10^-4=T
=> T=0.108 N